A grenade flying at a speed of 20 m / s was split into two pieces of the same mass. One of these pieces stays still after splitting. The speed of the other piece is: 

 A) 20 m / s 

 B) 10 m / s 

 C) 40m / s 

 D) 30m / s


1
Expert's answer
2021-03-10T17:19:55-0500
"p_i=p_f,""mu=m_1v_1+m_2v_2,""mu=m_2v_2,""v_2=\dfrac{mu}{m_2}=\dfrac{2m_2}{m_2}\cdot20\dfrac{m}{s}=40\dfrac{m}{s}."

Answer:

C) 40m / s